OpenStudy (anonymous):

solve the system by graphing {-2x=2y-4 {2x-y=-5 Help walk through this please.

OpenStudy (freckles):

-2x=2y-4 multiplying both sides of this equation by -1 gives 2x=-2y+4 so replace 2x in the second equation with (-2y+4)

OpenStudy (freckles):

oops solve by grahping

OpenStudy (freckles):

forgot to read that part

OpenStudy (freckles):

write both equations in the form y=mx+b this form is fun to play with because it tells you the slope m and the y-intercept b

OpenStudy (anonymous):

-2x=2y-4 2x=-y-5

OpenStudy (anonymous):

@freckles ?

OpenStudy (freckles):

yes you need to write those in y=mx+b form

OpenStudy (freckles):

for example -2x=2y-4 can be written in that form by doing the following: add 4 to both sides -2x+4=2y 2y=-2x+4 (just wrote the equation the other way around) divide 2 on both sides y=-x+2 so here you see we have a slope of -1 and a y-intercept of 2

OpenStudy (freckles):



OpenStudy (freckles):

try doing this with the other line you have there

OpenStudy (anonymous):

2x=-y-5 so do i add the 5 to 2x?

OpenStudy (anonymous):

here is a nice picture http://www.wolframalpha.com/input/?i=-2x%3D2y-4+%2C2x-y%3D-5

OpenStudy (anonymous):

oh okay